A projectile is thrown upward so that its distance above the ground after t seconds is given by the function h(t) = - 16t2 + 640t. After how many seconds does the projectile take to reach its maximum height? Show your work.
Answer by nerdybill(7384)

Given:
h(t) = -16t^2 + 640t
This is a "quadratic" (think parabola) and because the coefficient associated with the t^2 term is NEGATIVE, we know it opens downward.
Therefore, the vertex is the maximum.
.
The time (secs) when it reaches the max is:
t = -b/(2a)
t = -640/(2(-16))
t = -640/(-32)
t = 20 seconds